40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 10 01:10:56 2018, gautam, Update, LSC, Some vertex locking restored PRMIcarrier.png
    Reply  Tue Aug 14 00:27:55 2018, gautam, Update, LSC, Locking prep 
       Reply  Tue Aug 14 02:01:12 2018, gautam, Update, LSC, DRMI locking - partial success DRMI.png
Message ID: 14152     Entry time: Fri Aug 10 01:10:56 2018     Reply to this: 14160
Author: gautam 
Type: Update 
Category: LSC 
Subject: Some vertex locking restored 

For the first time after the whirlwind vent, I managed to lock the PRMI.

  • First, I did POX/POY locking, dither aligned the arms to maximize TRX and TRY.
  • Next, I misaligned the ETM and tested the Michelson locking
    • Since we've lost ~70% of power on the AS55 PD, I set the whitening gain for AS55 I and Q channels to +6dB (old value was 0dB).
    • worked alright. In this new config, the peak-to-peak Michelson fringe count is ~80 cts, while I reported ~60cts-pp a couple of months ago, so all seems good on that front.
    • But the config script in the IFOconfigure MEDM screen somehow doesn't set the AS55_Q ----> MICH_A element in the LSC input matrix anymore.
    • I edited the .snap file for this configuration to set the relevant matrix element EPICS channel to +1.0.
    • I also edited the overall loop gain for this configuration from +30 to +2 (for bright fringe, use -2 for dark fringe).
  • Feeling adventerous, I decided to try PRMI in the carrier resonant tuning (to be clear, PRCL on REFL11_I, MICH on AS55_Q).
    • Finding the REFL spot on the camera took a while since the PRM has been macroscopically misaligned for the mode-scanning
    • Went out to the table and centered the REFL beam onto REFL11 and REFL55 PDs - didn't need much tweaking, which is a good sign, since we shouldn't have screwed anything up on the symmetric side by any of the vent activities.
    • Restored PRMI locking using the IFOconfigure MEDM screen - lock caught almost immediately.
    • Ran the dither alignment servos for MICH and PRCL - BS needed a bit of encouragement to make the dark spot dark, but POP has been pretty stable over ~15mins.
    • I didn't take any loop transfer functions, to do.

I don't have the energy to make a DRMI attempt tonight - but the signs are encouraging. I'd like to use the IFO in the next few days to try and recover DRMI locking. The main concern is that the optical path on the AS beam has changed by ~0.3m I estimate. So the demod phase for AS55 may need to be adjusted, but the change due to optical path length only should be ~10degrees so the DRMI locking with the old settings should still work. Perhaps we also want to scan the PRC and SRC with the phase information from the Trans/Refl transfer functions as well.


Don't want to jinx it, but the c1lsc FE models have been stable. Tomorrow, I'd like to re-enable c1cal, since it has some useful channels for NBing. Could c1daf/c1oaf which have significant amounts of custom C code be the culprits?

Attachment 1: PRMIcarrier.png  70 kB  Uploaded Fri Aug 10 02:26:39 2018  | Hide | Hide all
PRMIcarrier.png
ELOG V3.1.3-